# DRAGON: Distributional Rewards Optimize Diffusion Generative Models

### v1

@@ B.3 Additional Aesthetics Score Optimization (RLHF) Analyses
As shown in Figure 7, when using DRAGON to optimize aesthetics score, the “middle-difficulty prompts”, i.e., prompts on which the baseline model achieves a near-average aesthetics score, see the most significant improvement. This makes sense, because highly rated music pieces are comparatively rare in the DMA dataset. As a result, as the aesthetics score improves throughout DRAGON training, the aesthetics predictor gradually becomes less reliable, and further improvements become more challenging. To understand this, notice in Figure 7(d) that after DRAGON training, the model generations’ aesthetics scores form a cluster whose most mass falls in the proximity of 0.9. As shown in Figure 6, the aesthetics model rarely makes predictions greater than 1 on its training dataset, and scores greater than 1.4 are almost never assigned. That is, DRAGON approaches the upper limit of the aesthetics model’s useful output range, where the aesthetics score becomes less reliable. Hence, if the aesthetics model can be improved in future work, we are confident that DRAGON optimization can further enhance music generation.
From a human preference alignment perspective, earlier image-domain approaches leveraged sparse human feedback (Pressman et al., 2022) or feedback on non-AI-generated examples that induce a distribution shift (Murray et al., 2012). [^2] While SAC now has over 238,000 images, it only had 4,000-5,000 when used to train the LAION aesthetics predictor. These weaker human preference signals are then augmented by training aesthetics predictors (Schuhmann & contributors, 2022) to create larger-scale synthetic preference datasets (Schuhmann, 2022). In some sense, DRAGON aesthetics score optimization follows a similar paradigm, but the creation of the synthetic dataset is now a part of the learning algorithm, and the demonstrations are on-policy. Hence, DRAGON sees less distribution shift and better utilizes sparse human preference. While more recent preference alignment methods leverage larger-scale human preference annotations for AI generation (Kirstain et al., 2023; Cideron et al., 2024), such data are highly expensive to collect, especially for the music modality where large-scale open-source preference datasets do not yet exist. We show that DRAGON addresses this challenge by learning from novel reward functions based on exemplar sets, achieving comparable results to directly learning from explicit human aesthetics ratings.
@@ B.4 Statistical Analyses on Human Listening Test Results
  • The posterior probability for DRAGON to outperform the baseline model (i.e. $>50\%$ win rate).
Since we collect binary human preferences, we model the observed win rate as a binomial distribution $K\sim\textrm{Binomial}(n,w)$, where $n=21\times 20=420$ is the total number of preference annotations, $w$ is DRAGON’s underlying true win rate which we aim to estimate, and the random variable $K$ is the number of positive ratings out of the $n$ data points. To evaluate the null hypothesis about DRAGON’s performance, we perform a binomial test using our observed number of wins $k_{\textrm{D}}$. We present the resulting $p$-values in Table 4. Intuitively, our results mean that if the true win rate $w$ were to be no greater than $50\%$, then the probabilities of obtaining our measured win rate would be $1.58\times 10^{-5}$ and $4.15\times 10^{-6}$ for the two DRAGON models. Since these are extremely small numbers, we can reject our null hypothesis with high confidence, concluding that DRAGON indeed outperforms the baseline.
@@
The closed-form solution is $w_{\textrm{lower-bound}}=\textrm{Beta}^{-1}(\alpha,k_{\textrm{D}},n-k_{\textrm{D}}+1)$, where $\textrm{Beta}^{-1}$ is the inverse cumulative distribution function of the beta distribution. Plugging in our observed $k_{\textrm{D}}$ values for the two DRAGON in this test, we obtain the $95\%$ confidence lower bound for $w$ shown in Table 4.
Finally, we use a Bayesian framework to analyze ${\mathbb{P}}(w>0.5\;|\;k_{\textrm{D}})$, the posterior probability for DRAGON’s true win rate $w$ to be greater than $50\%$. Using a non-informative uniform prior $\textrm{Uniform}(0,1)$, we have the prior density function $p(w)\propto 1$. Via Bayes’ Theorem, we have $p(w|k_{\textrm{D}})\propto p(k_{\textrm{D}}|w)\ p(w)$. Substituting $p(k_{\textrm{D}}|w)$ with the binomial mass function, we get $p(w|k_{\textrm{D}})\propto p^{k_{\textrm{D}}}(1-p)^{n-k_{\textrm{D}}}$, meaning that the posterior distribution is $\textrm{Beta}(k_{\textrm{D}}+1)(n-k_{\textrm{D}}+1)$. We can then use the cumulative distribution function of the beta distribution to compute ${\mathbb{P}}(w>0.5\;|\;k_{\textrm{D}})$. Table 4 shows that both DRAGON models have near-one posterior probability for $w>0.5$, and thus we say with high confidence that DRAGON outperforms its baseline.
Note that the above analysis framework implicitly assumes that all binary preferences are independent. However, this is not strictly satisfied because the ratings are grouped by raters. To model the individual preference of each rater, we consider the alternative approach of treating each rater’s observed win rate as a binomial variable. We then fit a generalized linear model (GLM) with a binomial family via logistic regression. Via the GLM approach, the $95\%$ confidence lower bound for $w$ is $56.25\%$ and $56.97\%$ for the two DRAGON models, and the $p$-values are $1.55\times 10^{-5}$ and $4.25\times 10^{-6}$. These numbers are extremely close to the simple binomial modeling results shown in Table 4, meaning that while the independence assumption is not strictly satisfied, the effect of this inaccuracy is minuscule for our analyses.
@@ C.1 Per-Song FAD Correlation Analysis
${\mathcal{D}}_{\textrm{ref}}$947 .106
${\mathcal{D}}_{\textrm{DRAGON-Aes}}$920 .116
${\mathcal{D}}_{+\textrm{persong}}$858 .092
${\mathcal{D}}_{+\textrm{dataset}}$903 .074
per-prompt SRCC reflects “the degree to which ranking per-song FAD agrees with ranking aesthetics score,” which is particularly meaningful for DRAGON training.
@@
As shown in Figure 8, the per-song FAD that uses ALIM audio as the reference statistics is noticeably correlated with human preference. Hence, it is a valid music quality measure, and we can expect optimizing it to enhance the model. Note that per-song FAD’s correlation with predicted aesthetics score is systematically more prominent than per-song FAD’s correlation with human-provided aesthetics annotations, especially when per-song FAD uses the same CLAP embeddings in the aesthetics model. This observation suggests that the aesthetics model predictions are less noisy than human-predicted ones, but may bring bias. Surprisingly, using ALIM’s textual music descriptions’ CLAP embeddings as the reference produces stronger correlations than using ALIM audio embeddings as the reference, implying that high-quality audio captions can offer rich and powerful insights into music quality. One explanation for text embeddings being more powerful than audio ones is that CLAP’s text encoder, RoBERTa (Liu et al., 2019), was pre-trained with datasets that covered much wider topics than CLAP’s audio encoder HTS-AT (Chen et al., 2022). Per-song text-FAD’s high statistical correlation to human preference explains the high performance of the DRAGON model that optimizes this quantity (as shown in Table 1(a)), which enhances content creation without expensive human ratings based on text-only data.
@@
Connection between per-song and full-dataset FAD. To understand the connection between per-song and full-dataset FAD, we analyze whether constructing ${\mathcal{D}}_{+}$ that optimizes one quantity also improves the other. We take the test generations from the baseline model and the DRAGON model that optimizes music aesthetics with the DPO loss, denoted as ${\mathcal{D}}_{\textrm{ref}}$ and ${\mathcal{D}}_{\textrm{DRAGON-Aes}}$ respectively, resulting in 2185 generation pairs. From each pair, we select the sample with a lower per-song FAD, and collect the 2185 chosen generations as ${\mathcal{D}}_{+\textrm{persong}}$. Meanwhile, we use Algorithm 1 to produce ${\mathcal{D}}_{+\textrm{dataset}}$, which also selects one piece from each of the 2185 pairs, but instead directly minimizes the full-dataset FAD.
@@
Table 5 presents the per-song and full-dataset FAD of ${\mathcal{D}}_{\textrm{ref}}$, ${\mathcal{D}}_{\textrm{DRAGON-Aes}}$, ${\mathcal{D}}_{+\textrm{persong}}$, and ${\mathcal{D}}_{+\textrm{dataset}}$. We observe that the dataset FAD of ${\mathcal{D}}_{+\textrm{persong}}$ is lower than that of both ${\mathcal{D}}_{\textrm{ref}}$ and ${\mathcal{D}}_{\textrm{DRAGON-Aes}}$, but higher than ${\mathcal{D}}_{+\textrm{dataset}}$. Conversely, the average per-song FAD of ${\mathcal{D}}_{+\textrm{dataset}}$ is lower than both ${\mathcal{D}}_{\textrm{ref}}$ and ${\mathcal{D}}_{\textrm{DRAGON-Aes}}$, but higher than ${\mathcal{D}}_{+\textrm{persong}}$. We can thus conclude that per-song FAD is correlated with full-dataset FAD, but if our goal is to minimize one of these two metrics, then directly optimizing that metric is more powerful than using the other as a surrogate. This observation highlights the importance of DRAGON’s unique capability to directly optimize reward functions that evaluate distributions, such as full-dataset FAD.
@@ C.2 MA Versus FADTK CLAP Embeddings for FAD Calculation
As discussed in Section 5.1, we consider two settings to compute CLAP embeddings for FAD calculation (for both per-song and full-dataset variants): MA and FADTK. The MA setting was determined via the ablation studies in Section B.2. It up-samples the 32-second 44.1kHz generations to 48kHz and uniformly splits them into eight partially overlapping 10-second audio chunks. The MA setting additionally down-samples the entire 32-second waveform to match the CLAP input sequence length and uses it to represent global information. Before encoding each chunk, we perform peak normalization to a range of [.5,.5].5.5[-.5,.5][ - .5 , .5 ], preventing reward hacking by merely manipulating the output magnitude. Eventually, the MA setting extracts nine 512-dimensional CLAP embeddings. The FADTK setting uses the same model checkpoint and up-sampling setting as MA. However, peak normalization is instead performed on the entire 32-second piece, which is subsequently divided into 10-second chunks with a 1-second hop length. As a result, FADTK produces more embeddings (25 instead of 9) per generation. The down-sampled chunk in the MA setting is not included.
For aesthetics score and CLAP score calculations, we obtain a score for each of the nine MA CLAP embeddings and average them. FAD (per-song and full-dataset) to text reference embeddings is also based on the MA embeddings due to their cross-modality nature shared with CLAP score. Meanwhile, since FAD to audio references is not cross-modal, we use the FADTK setting, which is reliable and standard in the literature. To compute Vendi score for a batch of music, we average the nine MA CLAP embeddings for each piece, and compute the Vendi score following eq. 4 with the per-song mean embeddings in this batch.
@@
In this section, we ablate the influence of these three differences by encoding human-created music with various encoding settings and computing the FAD between the embedding distributions resulting from different settings. These FAD numbers are listed in Figure 9(a). On the open-source SDNV dataset, we gradually modify the MA setting, removing the three main differences one by one to get closer to the FADTK setting, and examine how the FAD between the modified setting and the FADTK setting changes. The FAD between the unmodified MA setting and FADTK is almost as large as the FAD between two different audio datasets (SDNV-versus-ALIM) at $0.0688$. When we remove the down-sampled block from the MA setting, the FAD decreases to $0.0439$ but is still quite large. If we further remove peak normalization from both settings (now the only difference is hop size), then the FAD becomes only $0.0146$. Comparing these numbers, we conclude that all three differences between FADTK and MA contribute to their discrepancies, with peak normalization asserting the strongest influence and hop size mattering the least.
@@
In Figure 9(b), we compare the SDNV and ALIM datasets’ CLAP embedding distributions, considering FADTK and MA music embeddings as well as caption text embeddings. In both FADTK and MA settings, the FAD between SDNV and ALIM audio embeddings is about $0.1$. The text embedding FAD between the two datasets is much larger at $0.48$. In terms of audio-text correspondence, SDNV music descriptions are statistically closer to music pieces in the CLAP embedding space, with an FAD of around $0.65$. In comparison, ALIM captions see a large FAD of nearly $1$ from the corresponding audio data. For both ALIM and SDNV, MA audio embeddings are closer to the text embeddings. This observation supports using the MA setting when computing the FAD between generated audio and ground-truth text.
Next, we analyze how the differences between FADTK and MA affect their role as per-song FAD reference statistics for evaluating generated music. As shown in Figure 8, when ALIM or SDNV audio embeddings are used as reference (the setting to encode generated music matches with the reference setting for consistency), the per-song FAD computed with the MA setting is more strongly correlated with human preference. In contrast, the FADTK setting sees a much lower PLCC and a near-zero per-prompt SRCC with human ratings, suggesting that per-song FADTK makes decisions almost entirely based on semantic information, rather than aesthetics. If we remove peak normalization from FADTK, then we recover some correlation with human feedback, implying that per-song FADTK also attends to loudness. Gui et al. (2024) proposed to use per-song FADTK to predict music quality and identify outliers, and our result shows that such capabilities may be a result of the correlation between aesthetics, loudness, and semantics. This makes sense, because Gui et al. (2024) showed that the quality ratings given by per-song FADTK (based on audio) and large language models (GPT-4, based on captions) strongly agree with each other.

### v2

@@ B.3 Additional Aesthetics Score Optimization (RLHF) Analyses
As shown in Figure˜7, when using DRAGON to optimize aesthetics score, the “middle-difficulty prompts”, i.e., prompts on which the baseline model achieves a near-average aesthetics score, see the most significant improvement. This makes sense, because highly rated music pieces are comparatively rare in the DMA dataset. As a result, as the aesthetics score improves throughout DRAGON training, the aesthetics predictor gradually becomes less reliable, and further improvements become more challenging. To understand this, notice in Figure˜7(d) that after DRAGON training, the model generations’ aesthetics scores form a cluster whose most mass falls in the proximity of 0.9. As shown in Figure˜6, the aesthetics model rarely makes predictions greater than 1 on its training dataset, and scores greater than 1.4 are almost never assigned. That is, DRAGON approaches the upper limit of the aesthetics model’s useful output range, where the aesthetics score becomes less reliable. Hence, if the aesthetics model can be improved in future work, we are confident that DRAGON optimization can further enhance music generation. From a human preference alignment perspective, earlier image-domain approaches leveraged sparse human feedback (SAC) or feedback on non-AI-generated examples that induce a distribution shift (AVA). [^1] While SAC now has over 238,000 images, it only had 4,000-5,000 when used to train the LAION aesthetics predictor. These weaker human preference signals are then augmented by training aesthetics predictors (LAION-Aesthetics-Predictor-V2) to create larger-scale synthetic preference datasets (LAION-Aesthetics-Dataset). In some sense, DRAGON aesthetics score optimization follows a similar paradigm, but the creation of the synthetic dataset is now a part of the learning algorithm, and the demonstrations are on-policy. Hence, DRAGON sees less distribution shift and better utilizes sparse human preference. While more recent preference alignment methods leverage larger-scale human preference annotations for AI generation (Pick-A-Pic; MusicRL), such data are highly expensive to collect, especially for the music modality where large-scale open-source preference datasets do not yet exist. We show that DRAGON addresses this challenge by learning from novel reward functions based on exemplar sets, achieving comparable results to directly learning from explicit human aesthetics ratings.
@@ B.4 Statistical Analyses on Human Listening Test Results
  • The posterior probability for DRAGON to outperform the baseline model (i.e. $50\%$ win rate).
Since we collect binary human preferences, we model the observed win rate as a binomial distribution $K\sim\textrm{Binomial}(n,w)$, where $n=21\times 20=420$ is the total number of preference annotations, $w$ is DRAGON’s underlying true win rate which we aim to estimate, and the random variable $K$ is the number of positive ratings out of the $n$ data points. To evaluate the null hypothesis about DRAGON’s performance, we perform a binomial test using our observed number of wins $k_{\textrm{D}}$. We present the resulting $p$-values in Table˜5. Intuitively, our results mean that if the true win rate $w$ were to be no greater than $50\%$, then the probabilities of obtaining our measured win rate would be $1.58\times 10^{-5}$ and $4.15\times 10^{-6}$ for the two DRAGON models. Since these are extremely small numbers, we can reject our null hypothesis with high confidence, concluding that DRAGON indeed outperforms the baseline.
@@
The closed-form solution is $w_{\textrm{lower-bound}}=\textrm{Beta}^{-1}(\alpha,k_{\textrm{D}},n-k_{\textrm{D}}+1)$, where $\textrm{Beta}^{-1}$ is the inverse cumulative distribution function of the beta distribution. Plugging in our observed $k_{\textrm{D}}$ values for the two DRAGON in this test, we obtain the $95\%$ confidence lower bound for $w$ shown in Table˜5.
Finally, we use a Bayesian framework to analyze ${\mathbb{P}}(w>0.5\;|\;k_{\textrm{D}})$, the posterior probability for DRAGON’s true win rate $w$ to be greater than $50\%$. Using a non-informative uniform prior $\textrm{Uniform}(0,1)$, we have the prior density function $p(w)\propto 1$. Via Bayes’ Theorem, we have $p(w|k_{\textrm{D}})\propto p(k_{\textrm{D}}|w)\ p(w)$. Substituting $p(k_{\textrm{D}}|w)$ with the binomial mass function, we get $p(w|k_{\textrm{D}})\propto p^{k_{\textrm{D}}}(1-p)^{n-k_{\textrm{D}}}$, meaning that the posterior distribution is $\textrm{Beta}(k_{\textrm{D}}+1)(n-k_{\textrm{D}}+1)$. We can then use the cumulative distribution function of the beta distribution to compute ${\mathbb{P}}(w>0.5\;|\;k_{\textrm{D}})$. Table˜5 shows that both DRAGON models have near-one posterior probability for $w>0.5$, and thus we say with high confidence that DRAGON outperforms its baseline.
Note that the above analysis framework implicitly assumes that all binary preferences are independent. However, this is not strictly satisfied because the ratings are grouped by raters. To model the individual preference of each rater, we consider the alternative approach of treating each rater’s observed win rate as a binomial variable. We then fit a generalized linear model (GLM) with a binomial family via logistic regression. Via the GLM approach, the $95\%$ confidence lower bound for $w$ is $56.25\%$ and $56.97\%$ for the two DRAGON models, and the $p$-values are $1.55\times 10^{-5}$ and $4.25\times 10^{-6}$. These numbers are extremely close to the simple binomial modeling results shown in Table˜5, meaning that while the independence assumption is not strictly satisfied, the effect of this inaccuracy is minuscule for our analyses.
@@ C.1 Per-Song FAD Correlation Analysis
${\mathcal{D}}_{\textrm{ref}}$ .947 .106
${\mathcal{D}}_{\textrm{DRAGON-Aes}}$ .920 .116
${\mathcal{D}}_{+\textrm{persong}}$ .858 .092
${\mathcal{D}}_{+\textrm{dataset}}$ .903 .074
@@
per-prompt SRCC reflects “the degree to which ranking per-song FAD agrees with ranking aesthetics score,” which is particularly meaningful for DRAGON training. As shown in Figure˜8, the per-song FAD that uses ALIM audio as the reference statistics is noticeably correlated with human preference. Hence, it is a valid music quality measure, and we can expect optimizing it to enhance the model. Note that per-song FAD’s correlation with predicted aesthetics score is systematically more prominent than per-song FAD’s correlation with human-provided aesthetics annotations, especially when per-song FAD uses the same CLAP embeddings in the aesthetics model. This observation suggests that the aesthetics model predictions are less noisy than human-predicted ones, but may bring bias. Surprisingly, using ALIM’s textual music descriptions’ CLAP embeddings as the reference produces stronger correlations than using ALIM audio embeddings as the reference, implying that high-quality audio captions can offer rich and powerful insights into music quality. One explanation for text embeddings being more powerful than audio ones is that CLAP’s text encoder, RoBERTa (RoBERTa), was pre-trained with datasets that covered much wider topics than CLAP’s audio encoder HTS-AT (HTS-AT). Per-song text-FAD’s high statistical correlation to human preference explains the high performance of the DRAGON model that optimizes this quantity (as shown in Table˜2(a)), which enhances content creation without expensive human ratings based on text-only data.
@@
Connection between per-song and full-dataset FAD. To understand the connection between per-song and full-dataset FAD, we analyze whether constructing $\mathcal{D}$$_{+}$ that optimizes one quantity also improves the other. We take the test generations from the baseline model and the DRAGON model that optimizes music aesthetics with the DPO loss, denoted as ${\mathcal{D}}_{\textrm{ref}}$ and ${\mathcal{D}}_{\textrm{DRAGON-Aes}}$ respectively, resulting in 2185 generation pairs. From each pair, we select the sample with a lower per-song FAD, and collect the 2185 chosen generations as ${\mathcal{D}}_{+\textrm{persong}}$. Meanwhile, we use Algorithm˜1 to produce ${\mathcal{D}}_{+\textrm{dataset}}$, which also selects one piece from each of the 2185 pairs, but instead directly minimizes the full-dataset FAD.
@@
Table˜6 presents the per-song and full-dataset FAD of ${\mathcal{D}}_{\textrm{ref}}$, ${\mathcal{D}}_{\textrm{DRAGON-Aes}}$, ${\mathcal{D}}_{+\textrm{persong}}$, and ${\mathcal{D}}_{+\textrm{dataset}}$. We observe that the dataset FAD of ${\mathcal{D}}_{+\textrm{persong}}$ is lower than that of both ${\mathcal{D}}_{\textrm{ref}}$ and ${\mathcal{D}}_{\textrm{DRAGON-Aes}}$, but higher than ${\mathcal{D}}_{+\textrm{dataset}}$. Conversely, the average per-song FAD of ${\mathcal{D}}_{+\textrm{dataset}}$ is lower than both ${\mathcal{D}}_{\textrm{ref}}$ and ${\mathcal{D}}_{\textrm{DRAGON-Aes}}$, but higher than ${\mathcal{D}}_{+\textrm{persong}}$. We can thus conclude that per-song FAD is correlated with full-dataset FAD, but if our goal is to minimize one of these two metrics, then directly optimizing that metric is more powerful than using the other as a surrogate. This observation highlights the importance of DRAGON’s unique capability to directly optimize reward functions that evaluate distributions, such as full-dataset FAD.
@@ C.2 MA Versus FADTK CLAP Embeddings for FAD Calculation
As discussed in Section˜5.1, we consider two settings to compute CLAP embeddings for FAD calculation (for both per-song and full-dataset variants): MA and FADTK. The MA setting was determined via the ablation studies in Section˜B.2. It up-samples the 32-second 44.1kHz generations to 48kHz and uniformly splits them into eight partially overlapping 10-second audio chunks. The MA setting additionally down-samples the entire 32-second waveform to match the CLAP input sequence length and uses it to represent global information. Before encoding each chunk, we perform peak normalization to a range of [.5,.5][-.5,.5], preventing reward hacking by merely manipulating the output magnitude. Eventually, the MA setting extracts nine 512-dimensional CLAP embeddings. The FADTK setting uses the same model checkpoint and up-sampling setting as MA. However, peak normalization is instead performed on the entire 32-second piece, which is subsequently divided into 10-second chunks with a 1-second hop length. As a result, FADTK produces more embeddings (25 instead of 9) per generation. The down-sampled chunk in the MA setting is not included.
For aesthetics score and CLAP score calculations, we obtain a score for each of the nine MA CLAP embeddings and average them. FAD (per-song and full-dataset) to text reference embeddings is also based on the MA embeddings due to their cross-modality nature shared with CLAP score. Meanwhile, since FAD to audio references is not cross-modal, we use the FADTK setting, which is reliable and standard in the literature. To compute Vendi score for a batch of music, we average the nine MA CLAP embeddings for each piece, and compute the Vendi score following eq.˜4 with the per-song mean embeddings in this batch.
@@
In this section, we ablate the influence of these three differences by encoding human-created music with various encoding settings and computing the FAD between the embedding distributions resulting from different settings. These FAD numbers are listed in Figure˜9(a). On the open-source SDNV dataset, we gradually modify the MA setting, removing the three main differences one by one to get closer to the FADTK setting, and examine how the FAD between the modified setting and the FADTK setting changes. The FAD between the unmodified MA setting and FADTK is almost as large as the FAD between two different audio datasets (SDNV-versus-ALIM) at $0.0688$. When we remove the down-sampled block from the MA setting, the FAD decreases to $0.0439$ but is still quite large. If we further remove peak normalization from both settings (now the only difference is hop size), then the FAD becomes only $0.0146$. Comparing these numbers, we conclude that all three differences between FADTK and MA contribute to their discrepancies, with peak normalization asserting the strongest influence and hop size mattering the least.
@@
In Figure˜9(b), we compare the SDNV and ALIM datasets’ CLAP embedding distributions, considering FADTK and MA music embeddings as well as caption text embeddings. In both FADTK and MA settings, the FAD between SDNV and ALIM audio embeddings is about $0.1$. The text embedding FAD between the two datasets is much larger at $0.48$. In terms of audio-text correspondence, SDNV music descriptions are statistically closer to music pieces in the CLAP embedding space, with an FAD of around $0.65$. In comparison, ALIM captions see a large FAD of nearly $1$ from the corresponding audio data. For both ALIM and SDNV, MA audio embeddings are closer to the text embeddings. This observation supports using the MA setting when computing the FAD between generated audio and ground-truth text.
Next, we analyze how the differences between FADTK and MA affect their role as per-song FAD reference statistics for evaluating generated music. As shown in Figure˜8, when ALIM or SDNV audio embeddings are used as reference (the setting to encode generated music matches with the reference setting for consistency), the per-song FAD computed with the MA setting is more strongly correlated with human preference. In contrast, the FADTK setting sees a much lower PLCC and a near-zero per-prompt SRCC with human ratings, suggesting that per-song FADTK makes decisions almost entirely based on semantic information, rather than aesthetics. If we remove peak normalization from FADTK, then we recover some correlation with human feedback, implying that per-song FADTK also attends to loudness. FADTK proposed to use per-song FADTK to predict music quality and identify outliers, and our result shows that such capabilities may be a result of the correlation between aesthetics, loudness, and semantics. This makes sense, because FADTK showed that the quality ratings given by per-song FADTK (based on audio) and large language models (GPT-4, based on captions) strongly agree with each other.